# Physics-informed renormalisation group flows

### v2

@@ III.2 Functional optimisation of physics-informed renormalisation group flows
In this Section we only lay the foundation of generalised functional optimisation, further considerations and applications of functional optimisation in PIRG flows will be considered elsewhere. Hence, here we just briefly recapitulate functional optimisation as introduced in [6] and extended in [36], and then discuss the required generalisation. Functional optimisation is based on the minimisation of RG-trajectories. While this is a global criterion, in most practical applications local variants are used, partly because of the difficulty to define proper distances in the space of theories. The latter can be derived from the Fisher information metric, but its practical implementation is intricate. Still, minimising the RG-trajectory entails that at a given physical cutoff scale the remaining flow is minimal. In absence of flowing fields the flow operator has the representation
@@
${\cal O}[\phi,R_{k}]\in\left\{\hat{\cal O}\left[J,\frac{\delta}{\delta J}\right]Z_{k}[J]\right\}\,,$ (35)
where $\hat{\cal O}[x,y]$ is a general function of two variables, for more details see [6]. This representation can be readily obtained by a $\phi$-derivative of the flow of the effective action: The current itself is part of the set of operators $\{\cal O\}$, while the effective action is not.
@@
$\bigl{\|}G[\phi,0]-G[\phi,R^{\textrm{opt}}_{k}]\bigr{\|}\leq\bigl{\|}G[\phi]-G[\phi,R_{k}]\bigr{\|}\,,$ (37)
for all hyper surfaces $R_{k}\in\{R_{\bot}\}$ with the same physics. The latter condition is an intricate one and is related to the definition of a metric on the space of theories. A practical workaround is the definition of this hyper surface via the spectrum of the theory with a given cutoff. In short, we define $R_{k}\in\{R_{\bot}\}$ with
$\{R_{\bot}\}=\left\{R_{k}\,\Bigl{|}\,k^{2}_{\textrm{phys}}=\min_{\textrm{Re}\left(k_{\textrm{pol}}\right)}\{\left(\textrm{Re}\left(k_{\textrm{pol}}\right)\right)^{2}\}\right\}\,,$ (38a)
@@
$\{k_{\textrm{pol}}\}\ =\left\{k\,\Bigl{|}\,\Gamma^{(2)}[\phi,R_{k}](-k^{2}_{\textrm{pole}})=0\right\}\,.$ (38b)
@@ III.3.1 Effective action with fundamental and composite fields
$\Gamma[\varphi]=\Gamma\bigl{[}\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\bigr{]}\,,\quad\Delta\phi=(\phi_{2},...,\phi_{n})\,,$ (41)
and the $\phi^{a}_{\textrm{\tiny{EoM}}}$ are the solutions of the equation of motion (EoM) for $\phi_{a}$,
$\phi^{a}_{\textrm{\tiny{EoM}}}:\quad\left.\frac{\delta\Gamma}{\delta\phi_{a}}\right|_{\phi^{a}=\phi^{a}_{\textrm{\tiny{EoM}}}}=0\,.$ (42)
Equations 42 and 9 entail $J^{a}_{\phi}=0$ and the path integral 6 reduces to that with a source to the fundamental field, as does the Legendre transform 9,
$\Gamma\left[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\right]=\sup_{J_{\varphi}}\bigl{(}\int_{x}J_{\varphi}\varphi-\ln Z[J_{\varphi}]\bigr{)}\,.$ (43)
Equation 43 also provides a relation between the correlation functions of the fundamental field and that of the composite field as well as $\phi_{a}^{\textrm{EoM}}[\varphi]$. We elucidate this at the example of the inverse propagator of the fundamental field,
$\Gamma^{(2)}[\varphi](x,y)=$ $\,\left.\frac{\delta^{2}\Gamma[\varphi,\Delta\phi]}{\delta\varphi(x)\delta\varphi(y)}\right|_{\Delta\phi=\Delta\phi_{\textrm{\tiny{EoM}}}}$
$-\int_{\boldsymbol{x}_{a},\boldsymbol{x}_{b}}\Gamma_{ab}[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}]\frac{\delta\Delta\phi_{\textrm{\tiny{EoM}}}^{a}(\boldsymbol{x}_{a})}{\delta\varphi(x)}\frac{\delta\Delta\phi_{\textrm{\tiny{EoM}}}^{b}(\boldsymbol{x}_{b})}{\delta\varphi(y)}\,,$ (44)
@@
$\Gamma_{a_{1}\cdots a_{n}}[\phi]=\frac{\delta^{n}\Gamma[\phi]}{\delta\phi^{a_{1}}\cdots\delta\phi^{a_{n}}}\,.$ (45)
@@
$\frac{\delta^{n}}{\delta\varphi^{n}}\left[\left.\frac{\delta\Gamma}{\delta\Delta\phi_{a}}\bigl{[}\varphi,\Delta\phi[\varphi]\bigr{]}\right|_{\textrm{\tiny{EoM}}}f[\varphi]\right]\,,$ (46)
@@
$\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma}{\delta\Delta\phi_{a}}\right]\,,$ (47)
@@
$\frac{\delta^{2}\Gamma\left[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\right]}{\delta\varphi^{2}}=$ $\,\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma[\varphi,\Delta\phi]}{\delta\varphi}+\frac{\delta\Gamma}{\delta\Delta\phi^{a}}\frac{\delta\Delta\phi^{a}}{\delta\varphi}\right]$
$=\frac{\delta}{\delta\varphi}\frac{\delta\Gamma[\varphi,\Delta\phi]}{\delta\varphi}=\frac{\delta^{2}\Gamma[\varphi,\Delta\phi]}{\delta\varphi^{2}}+\frac{\delta\Delta\phi^{a}}{\delta\varphi}\frac{\delta^{2}\Gamma}{\delta\Delta\phi^{a}\delta\varphi}\,.$ (48)
@@
$\left.\frac{\delta^{2}\Gamma}{\delta\Delta\phi^{a}\delta\varphi}\right|_{\Delta\phi=\Delta\phi_{\textrm{\tiny{EoM}}}}$
$=\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma}{\delta\Delta\phi^{a}}[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}]\right]-\frac{\delta\Delta\phi^{b}}{\delta\varphi}\frac{\delta^{2}\Gamma[\varphi,\Delta\phi]}{\delta\Delta\phi^{b}\delta\Delta\phi^{a}}$
$=-\frac{\delta\Delta\phi^{b}}{\delta\varphi}\frac{\delta^{2}\Gamma[\varphi,\Delta\phi]}{\delta\Delta\phi^{b}\delta\Delta\phi^{a}}\,.$ (49)
@@ III.3.2 Effective action without fundamental fields and cumulants-preserving flows
A relevant example for an effective action approach without the fundamental fields is density functions theory. In QCD we can consider a formulation, where we also coupling a source to the quark density, $J_{n}\bar{q}(x)\gamma^{0}q(x)$, introduces the density field $n(x)$ in the Legendre transform with
@@
$\Gamma_{n}[n]=\Gamma_{\phi}\bigl{[}q^{\ }_{\textrm{EoM}}[n],\bar{q}^{\ }_{\textrm{EoM}}[n],n]\bigr{]}=\Gamma_{\phi}\bigl{[}0,0,n]\,.$ (52)
@@
Figure 2: Reconstruction of connected correlation functions (cumulants) using 67 in a zero-dimensional test-case, see Appendix B. The coordinate map $\dot{\phi}$ implements a classical target action flow, which is discussed in Section V.1. The solution can be computed by either taking $\mu_{\varphi}$-derivatives of the cumulants-preserving generating functional $Z^{(c)}(0)$ or by $J_{\varphi}$-derivatives of $Z_{\varphi}(J_{\varphi})$, for details see Appendix B.
@@
Both relations can be extended to $\varphi(x)\neq 0$, but we would like to keep this examples simple.
@@
Then, the integrated $n$th order connected correlations (or cumulants) of the mass operator $\hat{\varphi}^{2}(x)$ and the $\phi^{4}$-interaction operator $\hat{\varphi}^{4}(x)$ in 8, including mixed ones, can be obtained by $n$th order $-\mu_{\varphi}$ (or $-\lambda_{\varphi}$) derivatives of $-\Gamma[\phi^{\ }_{\textrm{EoM}}]$. For the present example of cumulants of $\hat{\varphi}^{2}(x)$ we obtain
@@ IV Feed-down flows
where $\dot{\omega}$ stands for the flows of $\omega^{\ }_{\Gamma_{T}}$ and for the functional equations that determine the expansion coefficients $\omega_{\dot{\phi}}$. Equation 71 breaks the feed-back loop as no higher order operator is present in the functional equations for the lower order ones.
@@ IV.1 Feed-down flows in the derivative expansion
$\Gamma[\phi]=$ $\,\sum_{n}\sum_{m,\boldsymbol{r}_{m}}\int_{x}\frac{1}{n!}\gamma^{\boldsymbol{a}_{m}}_{\boldsymbol{\mu}_{2n}}(\phi,\boldsymbol{r}_{m})\left[\prod_{i=1}^{m}\partial^{r_{i}}\phi^{a_{i}}\right]_{\boldsymbol{\mu}_{2n}}\,,$ (75a)
@@
The zeroth order flowing field, $\dot{\phi}^{a}_{0}$, has already been introduced and used in [2] with the notation $-\frac{1}{2}\eta(\rho)$ for $f_{0}(\rho)$. If contracting $\dot{\phi}^{a}_{1}$ with $\phi^{a}$, the five $f_{i}$ and $\bar{f}_{i}$ terms reduce to three terms with the factors $\phi\partial^{2}\phi,(\partial\phi)^{2}$ and $(\phi\partial\phi)^{2}$, the coefficients being $f_{1}+\phi^{2}\bar{f}_{1},f_{2}$ and $f_{3}+\phi^{2}\bar{f}_{2}$ respectively. However, if contracting $f^{a}_{1}$ with derivative terms, e.g. with $\partial^{2}\phi^{a}$, the terms stay apart. Accordingly, the five terms feed down differently to higher orders in the derivative expansion. Note also that in the first order of the derivative expansion of the effective action the five $O(\partial^{2})$-invariants used in 80 are related by partial integration, and typically one only considers the second and third invariant.
@@
We proceed with extending the above discussion to higher orders of the derivative expansion: Assume we have already constructed a feed-down flow for the $n$th order of the derivative expansion. The $n+1$st order term $\phi^{\prime}{}^{a}_{n+1}$ in the flowing field is given in 78b. Its potential feed-back contribution to lower order terms solely originates in 88 with $\phi^{\prime}{}^{a}_{1}\to\phi^{\prime}{}^{a}_{n+1}$. The latter has the form
@@
×$\hskip 28.45274pt\times\sum_{j=1}^{m}\left[q^{r_{m+1}}\prod_{i=1}^{m}\partial^{r_{i}}\phi_{b_{i}}\right]_{\boldsymbol{\mu}_{2n+2}}\,,$ (92a)
@@
A general $\dot{\phi}^{\prime}_{n+1}(q)$ leads to a feed-back to all lower orders: The powers of the derivatives $20\leq r_{1}+\cdots r_{m}\leq 2n+$ 2 determines the lowest order of the feed-back for each term in 92a. The lowest order is the $(q^{2})^{n+1}$-term in $z^{ab}(\phi,q^{2})$ in 85 and it contributes to the feed-down condition 87 for the effective potential. The next order contributes to the first order equations and leads to three constraints similar to 89. Evidently, it can be satisfied without restricting the generality of the flowing fields.
@@
Equation 92 expresses the fact that the contributions of the $n+1$st order terms in $\dot{\phi}$ do not contribute to the orders $i<n+1$, the sufficient condition for feed-down flows. Note, that seemingly this may not allow us to satisfy all constraints for the flowing fields required to generate a given target action effective action. However, as in the first order, the derivative terms in a general expansion of $\dot{\phi}$, that have to be dropped due to 92, are linked to other terms still allowed by 92, if inserted in the effective action. For formulating this property we remark that a general $\dot{\phi}$ simply is a general functional in the given order of the derivative expansion, and hence all functionals can be described by it. In particular we have

### v3

@@ III.2 Functional optimisation of physics-informed renormalisation group flows
In this Section we only lay the foundation of generalised functional optimisation, further considerations and applications of functional optimisation in PIRG flows will be considered elsewhere. In particular, we have completed the PIRG approach with general operators flows in [40]. Here we just briefly recapitulate functional optimisation based on operator flows as introduced in [6] and extended in [41], and discuss the required generalisation. Functional optimisation is based on the minimisation of RG-trajectories. While this is a global criterion, in most practical applications local variants are used, partly because of the difficulty to define proper distances in the space of theories. The latter can be derived from the Fisher information metric, but its practical implementation is intricate. Still, minimising the RG-trajectory entails that at a given physical cutoff scale the remaining flow is minimal. In absence of flowing fields the flow operator has the representation
@@
${\cal O}[\phi,R_{\phi}]\in\left\{\hat{\cal O}\left[J,\frac{\delta}{\delta J}\right]Z_{k}[J]\right\}\,$, (35)
where $\hat{\cal O}[x,y]$ is a general function of two variables, for more details see [6]. For a specific class of flowing fields, 34 admits a similar structure, see [40]. Therefore, the arguments below are readily extended to PIRGs. This representation can be readily obtained by a $\phi$-derivative of the flow of the effective action: The current itself is part of the set of operators $\{\cal O\}$, while the effective action is not.
@@
$\bigl{\|}G[\phi,0]-G[\phi,R^{\textrm{opt}}_{k}]\bigr{\|}\leq\bigl{\|}G[\phi]-G[\phi,R_{\phi}]\bigr{\|}\,$, (37)
for all hyper surfaces $R_{\phi}\in\{R_{\bot}\}$ with the same physics. The latter condition is an intricate one and is related to the definition of a metric on the space of theories. A practical workaround is the definition of this hyper surface via the spectrum of the theory with a given cutoff. In short, we define $R_{\phi}\in\{R_{\bot}\}$ with
$\{R_{\bot}\}=\left\{R_{\phi}\,\Bigl{|}\,k^{2}_{\textrm{phys}}=\min_{\textrm{Re}\left(k_{\textrm{pol}}\right)}\{\left(\textrm{Re}\left(k_{\textrm{pol}}\right)\right)^{2}\}\right\}\,$, (38a)
@@
$\{k_{\textrm{pol}}\}\ =\left\{k\,\Bigl{|}\,\Gamma^{(2)}[\phi,R_{\phi}](-k^{2}_{\textrm{pole}})=0\right\}\,.$ (38b)
@@ III.3.1 Effective action with fundamental and composite fields
$\Gamma_{\varphi}[\varphi]=\Gamma_{\phi}\bigl{[}\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\bigr{]}\,,\quad\Delta\phi=(\phi_{2},...,\phi_{n})\,$, (41)
and the $\phi^{a}_{\textrm{\tiny{EoM}}}$ are the solutions of the equation of motion (EoM) for $\phi^{a}$,
$\phi^{a}_{\textrm{\tiny{EoM}}}:\quad\left.\frac{\delta\Gamma_{\phi}}{\delta\phi_{a}}\right|_{\phi^{a}=\phi^{a}_{\textrm{\tiny{EoM}}}}=0\,.$ (42)
Note, that we have taken the Legendre transform with respect to all fields $\phi=(\varphi,\Delta\phi)$. Equations 42 and 9 entails $J^{a}_{\phi}=0$ and the path integral 6 reduces to that with a source to the fundamental field, as does the Legendre transform 9,
$\Gamma_{\phi}\bigl{[}\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\bigr{]}=\sup_{J_{\varphi}}\bigl{(}\int_{x}J_{\varphi}\,\varphi-\ln Z_{\varphi}[J_{\varphi}]\bigr{)}\,$, (43)
with $Z_{\varphi}[J_{\varphi}]=Z_{\phi}[J_{\varphi},J_{\Delta\phi}=0]$. Equation 43 also provides a relation between the correlation functions of the fundamental field and that of the composite field as well as $\phi_{a}^{\textrm{EoM}}[\varphi]$. We elucidate this at the example of the inverse propagator of the fundamental field,
$\Gamma_{\varphi}^{(2)}[\varphi](x,y)=$ $\,\left.\frac{\delta^{2}\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\varphi(x)\delta\varphi(y)}\right|_{\Delta\phi=\Delta\phi_{\textrm{\tiny{EoM}}}}$
$\hskip-56.9055pt-\int\limits_{\boldsymbol{x}_{a},\boldsymbol{x}_{b}}\Gamma_{\phi,ab}[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}]\frac{\delta\Delta\phi_{\textrm{\tiny{EoM}}}^{a}(\boldsymbol{x}_{a})}{\delta\varphi(x)}\frac{\delta\Delta\phi_{\textrm{\tiny{EoM}}}^{b}(\boldsymbol{x}_{b})}{\delta\varphi(y)}\,$, (44)
@@
$\Gamma_{\phi,a_{1}\cdots a_{n}}[\phi]=\frac{\delta^{n}\Gamma_{\phi}[\phi]}{\delta\phi^{a_{1}}\cdots\delta\phi^{a_{n}}}\,.$ (45)
@@
$\frac{\delta^{n}}{\delta\varphi^{n}}\left[\left.\frac{\delta\Gamma_{\phi}}{\delta\Delta\phi_{a}}\bigl{[}\varphi,\Delta\phi[\varphi]\bigr{]}\right|_{\textrm{\tiny{EoM}}}f[\varphi]\right]\,$, (46)
@@
$\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma_{\phi}}{\delta\Delta\phi_{a}}\right]\,$, (47)
@@
$\frac{\delta^{2}\Gamma_{\phi}\left[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}[\varphi]\right]}{\delta\varphi^{2}}=$ $\,\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\varphi}+\frac{\delta\Gamma_{\phi}}{\delta\Delta\phi^{a}}\frac{\delta\Delta\phi^{a}}{\delta\varphi}\right]$
$\hskip-99.58464pt=\frac{\delta}{\delta\varphi}\frac{\delta\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\varphi}=\frac{\delta^{2}\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\varphi^{2}}+\frac{\delta\Delta\phi^{a}}{\delta\varphi}\frac{\delta^{2}\Gamma_{\phi}}{\delta\Delta\phi^{a}\delta\varphi}\,.$ (48)
@@
$\left.\frac{\delta^{2}\Gamma_{\phi}}{\delta\Delta\phi^{a}\delta\varphi}\right|_{\Delta\phi=\Delta\phi_{\textrm{\tiny{EoM}}}}$
$\hskip-71.13188pt=\frac{\delta}{\delta\varphi}\left[\frac{\delta\Gamma_{\phi}}{\delta\Delta\phi^{a}}[\varphi,\Delta\phi_{\textrm{\tiny{EoM}}}]\right]-\frac{\delta\Delta\phi^{b}}{\delta\varphi}\frac{\delta^{2}\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\Delta\phi^{b}\delta\Delta\phi^{a}}$
$\hskip-71.13188pt=-\frac{\delta\Delta\phi^{b}}{\delta\varphi}\frac{\delta^{2}\Gamma_{\phi}[\varphi,\Delta\phi]}{\delta\Delta\phi^{b}\delta\Delta\phi^{a}}\,.$ (49)
@@ III.3.2 Effective action without fundamental fields and cumulants-preserving flows
A relevant example for an effective action approach without the fundamental fields is density functions theory. For example, we can consider a formulation in QCD where a source is coupled to the quark density, $J_{n}\bar{q}(x)\gamma^{0}q(x)$. This introduces the density field $n(x)$ in the Legendre transform with
@@
$\Gamma_{n}[n]=\Gamma_{\phi}\bigl{[}q^{\ }_{\textrm{EoM}}[n],\bar{q}^{\ }_{\textrm{EoM}}[n],n\bigr{]}=\Gamma_{\phi}\bigl{[}0,0,n]\,.$ (52)
@@
Figure 2: Reconstruction cumulants using 67 in a zero-dimensional test-case, see Appendix B. The coordinate map $\dot{\phi}$ implements a classical target action flow, which is discussed in Section V.1. The solution can be computed by either taking $\mu_{\varphi}$-derivatives of the cumulants-preserving generating functional $Z^{(c)}(0)$ or by $J_{\varphi}$-derivatives of $Z_{\varphi}(J_{\varphi})$, for details see Appendix B.
@@
Both relations can be extended to $\varphi(x)\neq 0$, but we would like to keep these examples simple.
@@
Then, the integrated $n$th order or cumulants of the mass operator $\hat{\varphi}^{2}(x)$ and the $\phi^{4}$-interaction operator $\hat{\varphi}^{4}(x)$ in 8, including mixed ones, can be obtained by $n$th order $\mu_{\varphi}$ (or $\lambda_{\varphi}$) derivatives of $\Gamma[\phi^{\ }_{\textrm{EoM}}]$. For the present example of cumulants of $\hat{\varphi}^{2}(x)$ we obtain
@@ IV Feed-down flows
where $\dot{\omega}$ stands for the flows of $\omega^{\ }_{\Gamma_{T}}$ and for the functional equations that determine the expansion coefficients $\omega_{\dot{\phi}}$.
We visualise the feed-down structure of the flow in Figure 3. Equation 71 breaks the feed-back loop as no higher order operator is present in the functional equations for the lower order ones (dashed red arrow).
$(\Gamma_{T},\dot{\phi})_{n}\hskip 5.69054pt$$(\Gamma_{T},\dot{\phi})_{n+1}$
Figure 3: Feed-down structure of the flow with the feed-down condition 71. The straight back arrow indicates the feed-down of the $n$th-order pair into the higher order flows, specifically the n+1n+1st order. The dashed red arrow indicates the potential feed-back of the higher order terms into the lower order ones. The feed-down condition 71 breaks up this feed-back.
@@ IV.1 Feed-down flows in the derivative expansion
$\Gamma[\phi]=$ $\,\sum_{n}\sum_{m,\boldsymbol{r}_{m}}\int_{x}\frac{1}{n!}\gamma^{\boldsymbol{a}_{m}}_{\boldsymbol{\mu}_{2n}}(\phi,\boldsymbol{r}_{m})\left[\prod_{i=1}^{m}\partial^{r_{i}}\phi^{a_{i}}\right]_{\boldsymbol{\mu}_{2n}}\hskip-8.5359pt\,$, (75a)
@@
The zeroth order flowing field, $\dot{\phi}^{a}_{0}$, has already been introduced and used in [2] with the notation $\frac{1}{2}\eta(\rho)$ for $f_{0}(\rho)$. If contracting $\dot{\phi}^{a}_{1}$ with $\phi^{a}$, the five $f_{i}$ and $\bar{f}_{i}$ terms reduce to three terms with the factors $\phi\partial^{2}\phi,(\partial\phi)^{2}$ and $(\phi\partial\phi)^{2}$, the coefficients being $f_{1}+\phi^{2}\bar{f}_{1},f_{2}$ and $f_{3}+\phi^{2}\bar{f}_{2}$ respectively. However, if contracting $f^{a}_{1}$ with derivative terms, e.g. with $\partial^{2}\phi^{a}$, the terms stay apart. Accordingly, the five terms feed down differently to higher orders in the derivative expansion. Note also that in the first order of the derivative expansion of the effective action the five $O(\partial^{2})$-invariants used in 80 are related by partial integration, and typically one only considers the second and third invariant.
@@
We proceed with extending the above discussion to higher orders of the derivative expansion: Assume we have already constructed a feed-down flow for the $n$th order of the derivative expansion. The n+1n+1st order term $\phi^{\prime}{}^{a}_{n+1}$ in the flowing field is given in 78b. Its potential feed-back contribution to lower order terms solely originates in 88 with $\phi^{\prime}{}^{a}_{1}\to\phi^{\prime}{}^{a}_{n+1}$. The latter has the form
@@
$\hskip 28.45274pt\times\sum_{j=1}^{m}\left[q^{r_{m+1}}\prod_{i=1}^{m}\partial^{r_{i}}\phi_{b_{i}}\right]_{\boldsymbol{\mu}_{2n+2}}\hskip-8.5359pt\,$, (92a)
@@
A general $\dot{\phi}^{\prime}_{n+1}(q)$ leads to a feed-back to all lower orders: The powers of the derivatives $0\leq r_{1}+\cdots r_{m}\leq 2n+2$ determines the lowest order of the feed-back for each term in 92a. The lowest order is the $(q^{2})^{n+1}$-term in $z^{ab}(\phi,q^{2})$ in 85 and it contributes to the feed-down condition 87 for the effective potential. The next order contributes to the first order equations and leads to three constraints similar to 89. Evidently, it can be satisfied without restricting the generality of the flowing fields.
@@
Equation 92 expresses the fact that the contributions of the n+1n+1st order terms in $\dot{\phi}$ do not contribute to the orders $i<n+1$, the sufficient condition for feed-down flows. Note, that seemingly this may not allow us to satisfy all constraints for the flowing fields required to generate a given target action effective action. However, as in the first order, the derivative terms in a general expansion of $\dot{\phi}$, that have to be dropped due to 92, are linked to other terms still allowed by 92, if inserted in the effective action. For formulating this property we remark that a general $\dot{\phi}$ simply is a general functional in the given order of the derivative expansion, and hence all functionals can be described by it. In particular we have